OpenStudy (anonymous):

The molecular weight of CaCl2 is 110.98 g/mol. In order to make up 500. mL of a 0.500 M solution, you need: a)55.6 g b)27.7 g c)222.2 g d)111.1 g e)34.4 g

OpenStudy (jfraser):

Concentration in terms of M is (moles/L). When a solution is made, the moles of solute and total volume of the solution determine its concentration. Since \[C = mol/L\] then \[mol = C * L\] the moles of solute will be the product of the concentration and volume.\[mol = (0.5M)*(0.5L) = 0.25mol solute\] so 0.25moles of CaCl2 will have to be dissolved into 500mL (0.5L) of solution. Use the molar mass of CaCl2 to find the grams needed

OpenStudy (anonymous):

So which is it?